# A Taxonomy of Questions for Critical Reflection in Machine-Assisted Decision-Making

### v2

@@ 5. Evaluation of Our Taxonomy
To evaluate the proposed question taxonomy, we presented a first version with sample questions and the final version in two phases to the same orthopaedic surgeon and orthopaedic researcher in individual meetings. Their feedback was positive and they recognised the relevance of the taxonomy. As one remarked, given that many question types from the taxonomy are derived from data, it is difficult to ask about data points that are not contained in the dataset.111In this case, general questions like “Were all relevant information considered?” could help. Nevertheless, they were confident that questions could help to re-evaluate initial assumptions of the physician, and counteract both priming by information from a DSS and habitual priming, e.g., seeing similar yet different patient cases throughout the day. Moreover, they mentioned that the questions derived from the taxonomy could help to treat different patient cases more objectively by reducing discrepancies between these cases, such as different patients describing their pain differently, i.e., pain catastrophising.
5.1. Transferability
@@
To assess the applicability of our taxonomy in another domain, we discussed the finalised version with two postdoctoral researchers in the field of education and technology, and one doctoral researcher who is also a primary school teacher, in three individual feedback sessions of 30 minutes each. They all shared examples of so-called learning analytics dashboards (LAD), with which educators can plan lessons, define learning objectives, track the learning progress of students and create exercises accordingly. These LADs also allow teachers to use exercises and learning objectives that the system recommends based on the student’s performance. All three researchers therefore confirmed the expected relevance and applicability of our taxonomy for stimulating teachers’ critical reflection in interpreting the student’s progress and integrating recommendations of LADs. As one researcher commented: “Teachers feel the need to be more critical, but do not know how”, especially as these LADs can only take part of the student’s trajectory into account while informing and shaping decisions about the student’s future.
@@
After the three feedback sessions, we created sample questions for the educational domain (Table A4). We presented these questions, including our taxonomy, in a lab session attended by 12 people, from doctoral researchers to full professors. The research group, which explores the impacts of technologies on learning and teaching practices, recognised the usefulness of our taxonomy. As one participant commented, the sample questions we created are very close to the questions teachers already ask themselves. Nevertheless, the same researcher continued, our taxonomy could be informative to see which questions are covered and which are left out. In addition, another researcher noted that it is very promising that our taxonomy aligns with the questions that teachers ask in relation to algorithmic learning goal recommendations.
@@
In comparison to the medical sample questions (Table A3), the wording of the educational sample questions differs. The structure and elements of critical reflection however remain the same for both. We therefore maintain that our taxonomy is transferable and can be applied to other decision-making domains without adaptation as it builds on general concepts of critical thinking.
@@ 5.2. Effectiveness of Questions
Another type of evaluation, beyond the scope of this paper, is to assess the effectiveness of the questions created from the taxonomy in decision-making, and the extent to which they promote critical reflection and cognitive engagement of the decision-maker. Existing studies, primarily from the educational domain, point to the positive effects of Socratic questioning strategies on moral reasoning in nursing (Torabizadeh et al., 2018) and critical thinking skills during (healthcare) education of students (Ishtiaq et al., 2024; Jameel et al., 2025; Yang et al., 2005; Makhene, 2019; Ho et al., 2023). These studies demonstrate the validity and robustness of the Socratic questioning approach on which our proposed taxonomy is based, which makes the prospects of our taxonomy seem promising.
To measure reflection and cognitive engagement, it is important to consider the entire decision-making process, including the reasoning process, whether alternatives were considered, and how confident the decision-maker is in the final decision. Current evaluation methods for decision-making and overreliance on DSS, however, typically only focus on the outcome of a decision, i.e., decision accuracy (Guo et al., 2024; Klingbeil et al., 2024). As such, evaluating the effects of questions in promoting critical reflection and decision-making remains an open challenge. Existing methods for measuring cognitive engagement are associated with considerable effort and technical barriers, such as electroencephalography (EEG) (Hassib et al., 2017; Berka et al., 2007), or the measurement of pupil dilation (Van Der Wel and Van Steenbergen, 2018). Consequently, an evaluation framework would need to be developed in future work, e.g., a self-report scale that takes the entire decision-making process into account. For this, inspiration can be drawn from existing scales for factors associated with cognitive engagement, such as self-efficacy, motivation, and confidence (Lee et al., 2025; Hart et al., 2014; Greene, 2015), as well as from methods in the educational domain for assessing student’s cognitive engagement with technology (Vongkulluksn et al., 2022).
@@ 6. Discussion
We have presented a taxonomy of question types to support the systematic creation of questions that stimulate critical reflection during machine-assisted decision-making. To do so, we combined and utilised a taxonomy of Socratic questions and a question bank for explainable AI. We do not claim that there is a direct mapping between the taxonomy of Socratic questions and the question categories from the XAI question bank. Nevertheless, as Figure 1 shows, the questions identified for the human-centred XAI approach are very model- and data-centric. In that regard, and referring back to the need for explanations beyond the model (section 2.2), our taxonomy can help to take into account human factors like assumptions of the decision-maker and stakeholder preferences, as well as consequences of a decision. In doing so, questions can stimulate discussion about the machine recommendation and how it can be integrated into the broader decision-making process.
@@ 6.1. Different Stakeholders
Our taxonomy is primarily targeted at decision-makers who operate DSS to make informed decisions. Consequently, the aim of our taxonomy is not to formulate questions for the decision-maker who then evaluates the performance of the DSS. This would unfairly shift the responsibility from developers and regulators to operators. The deployed models must be evaluated and tested beforehand, especially in crucial settings such as healthcare, which requires a shift from individual human oversight to institutional oversight (Green, 2022). Considering this, other stakeholders could also benefit from critical reflection and thus our taxonomy. Developers, for example, might ask themselves which taken for granted assumptions they have while constructing a model and whether the used dataset is adequate to solve a given problem. Auditors assessing compliance could benefit from our taxonomy when defining objectives for an audit of the model or dataset (Raji et al., 2020; Birhane et al., 2024). Moreover, the taxonomy could be used as a form of supporting AI literacy to help people remain more critical of machine-generated information, especially in view of misinformation (Zavolokina et al., 2024). Patients could be supported in identifying (treatment) preferences, and asking relevant questions to their healthcare providers (operating DSS), or while interacting with a medical chatbot themselves. Students could be taught to question the relevance of retrieved information and could be helped to remain cognitively engaged in their technology-aided learning. Our proposed taxonomy could therefore benefit various stakeholders involved in the design, development, testing, and use of DSS or machine learning systems.
@@ 6.2. Creation of Questions
For questions to be meaningful and to support reflection effectively, researchers or designers ideally create them with input and feedback from domain experts, e.g., physicians. While creating the sample questions for the educational domain (Table A4) we realised that it is difficult to create questions without having much knowledge about the respective use case, access to data, and the functionality of the DSS or learning analytics dashboard (LAD). Nevertheless, uncovering these blind spots might enable further inquiry into how the system works, and open up a dialogue between developers or service providers and domain experts or affected organisations, such as hospitals or schools.
Furthermore, certain questions might fall into multiple categories of our taxonomy, as different question types can overlap or build on other categories, such as the causal structure of a recommendation (Q4) that can entail questions about the data (Q1&Q3). We believe, however, that this is not a major problem and could rather help to formulate questions more precisely so that they can be delineated. The overlap also shows how different elements or categories are connected to each other.
To help formulate questions, we have presented Bloom’s taxonomy. We argue that questions should address the higher levels of critical thinking, that is evaluating, analysing, and creating. As such, it might be that questions cannot be answered directly and lead to further questions (Paul, 2019). To motivate further inquiry and engagement of the decision-maker, it could be helpful, as mentioned by the orthopaedic surgeon we spoke to, to provide a brief explanation why a question should be entertained. For example, the statistical effect of the patient’s age on the outcome, according to the DSS, could be indicated in the question about the age. In view of this, however, it must also be considered that questions can be leading and point in certain directions, which might lead the decision-maker to make an undesirable decision. It is thus again important that domain experts are involved in the formulation of questions.
@@ 6.3. Presentation of Questions
Regarding the presentation of questions to decision-makers, one of the educational researchers we spoke to suggested that a poster of our taxonomy with sample questions could be displayed in the staff room to sensitise teachers to the different elements of the machine-assisted decision-making process. Alongside an analogue form and building on the above-mentioned proposals of DSS supporting deliberation and reflection (section 2.3), our taxonomy could inform the design of a technical implementation. As such, we envision a reflection machine (RM) that functions as an additional system or component of a DSS. The RM can take into account the input data (e.g., patient case information), the functioning of the DSS, e.g., through XAI methods such as feature importance, as well as the recommendations. With this information and the use of our taxonomy relevant questions can be generated and presented to the decision-maker. We advise against the use of language models, as the generated questions could be irrelevant and unreliable (Weidinger et al., 2022) and chatbots might lead to further overreliance (Anderl et al., 2024; Maeda and Quan-Haase, 2024). Besides, supporting the decision-making process with questions might be an unusual and unlearned interaction. Hence, evaluating the acceptance and usability (Lund, 2001; Venkatesh et al., 2016) of such an RM with specific questions is another consideration while assessing the effectiveness of questions.
@@ 6.4. Cognitive Load
While future work needs to evaluate the effectiveness of questions, it could be that questions interrupt the workflow and increase the cognitive load of the decision-maker. Similarly, in the aforementioned study, where explanations were presented at different times as a form of cognitive intervention, participants reported an increased cognitive load from these interventions, which is why they favoured them less (Buçinca et al., 2021). Ultimately, the cognitive burden, however, is unavoidable, and inherent to critical reflection. As John Dewey (1933) mentions on the difficulty of reflective thinking:
[…] yet thinking need not be reflective. For the person may not be sufficiently critical about the ideas that occur to him. He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled. One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching (p.17).
The aim of increasing cognitive engagement therefore leads to an increase in cognitive load. Nevertheless, the right amount of effort required and effectiveness needs to be found. Questions must be well integrated into decision-making processes in order to enable the operator to constructively take the posed question into account. This also depends at least in part on the competences, decision-making style, and need for cognition of the decision-maker.
@@ 7. Conclusion
We have provided a taxonomy of 10 question types aimed at promoting critical reflection in machine-assisted decision-making. Careful selection and presentation of these question types in the right circumstances (e.g., when time for reflection is available) can increase cognitive engagement and support the effective human oversight required by the European AI Act.
@@ Acknowledgements.
We thank Sebastiaan van de Groes and Miranda van Hooff, both at Radboud UMC, for their input on the construction of the taxonomy. Many thanks to Josh Ring and the Adaptive Learning Lab, in particular to Susan Verbunt-Janssen, Susanne de Mooij, and Zowi Vermeire for their valuable insights, and helping with the sample questions for the educational domain. Thanks to Cătălina Rățală and the Decision Science Group for their questions and feedback on a presentation of our paper. Comments from three anonymous ACM FAccT reviewers on an earlier version helped us to improve the paper. This research is funded by the Donders Centre for Cognition.References
@@
- Anderl et al. (2024) Christine Anderl, Stefanie H. Klein, Büsra Sarigül, Frank M. Schneider, Junyi Han, Paul L. Fiedler, and Sonja Utz. 2024. Conversational Presentation Mode Increases Credibility Judgements during Information Search with ChatGPT. Scientific Reports 14, 1 (July 2024), 17127. https://doi.org/10.1038/s41598-024-67829-6
@@
- Ely (2000) J. W Ely. 2000. A Taxonomy of Generic Clinical Questions: Classification Study. BMJ 321, 7258 (August 2000), 429–432. https://doi.org/10.1136/bmj.321.7258.429
@@
- Facione (2000) Peter A. Facione. 2000. The Disposition Toward Critical Thinking: Its Character, Measurement, and Relationship to Critical Thinking Skill. Informal Logic 20, 1 (January 2000). https://doi.org/10.22329/il.v20i1.2254
@@
- Hong and Choi (2011) Yi-Chun Hong and Ikseon Choi. 2011. Three Dimensions of Reflective Thinking in Solving Design Problems: A Conceptual Model. Educational Technology Research and Development 59, 5 (April 2011), 687–710. https://doi.org/10.1007/s11423-011-9202-9
@@
- Ma et al. (2024a) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma. 2024a. Towards Human-AI Deliberation: Design and Evaluation of LLM-Empowered Deliberative AI for AI-Assisted Decision-Making. arXiv:2403.16812 [cs]
- Ma et al. (2024b) Shuai Ma, Chenyi Zhang, Xinru Wang, Xiaojuan Ma, and Ming Yin. 2024b. Beyond Recommender: An Exploratory Study of the Effects of Different AI Roles in AI-Assisted Decision Making. arXiv:2403.01791 [cs]
- Maeda and Quan-Haase (2024) Takuya Maeda and Anabel Quan-Haase. 2024. When Human-AI Interactions Become Parasocial: Agency and Anthropomorphism in Affective Design. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. ACM, Rio de Janeiro Brazil, 1068–1077. https://doi.org/10.1145/3630106.3658956
@@
- Miller (2023) Tim Miller. 2023. Explainable AI Is Dead, Long Live Explainable AI! Hypothesis-driven Decision Support. arXiv:2302.12389 [cs]
@@
- Sandars (2009) John Sandars. 2009. The Use of Reflection in Medical Education: AMEE Guide No. 44. Medical Teacher 31, 8 (January 2009), 685–695. https://doi.org/10.1080/01421590903050374
@@
- Sengers et al. (2005) Phoebe Sengers, Kirsten Boehner, Shay David, and Joseph ’Jofish’ Kaye. 2005. Reflective Design. In Proceedings of the 4th Decennial Conference on Critical Computing: Between Sense and Sensibility. ACM, Aarhus Denmark, 49–58. https://doi.org/10.1145/1094562.1094569
@@
- van der Stigchel et al. (2023) Birgit van der Stigchel, Karel van den Bosch, Jurriaan van Diggelen, and Pim Haselager. 2023. Intelligent Decision Support in Medical Triage: Are People Robust to Biased Advice? Journal of Public Health 45, 3 (March 2023), 689–696. https://doi.org/10.1093/pubmed/fdad005
- van der Waa et al. (2021) Jasper van der Waa, Sabine Verdult, Karel van den Bosch, Jurriaan van Diggelen, Tjalling Haije, Birgit van der Stigchel, and Ioana Cocu. 2021. Moral Decision Making in Human-Agent Teams: Human Control and the Role of Explanations. Frontiers in Robotics and AI 8 (May 2021), 640647. https://doi.org/10.3389/frobt.2021.640647
@@
- van Hooff et al. (2014) Miranda L. van Hooff, Jan van Loon, Jacques van Limbeek, and Marinus de Kleuver. 2014. The Nijmegen Decision Tool for Chronic Low Back Pain. Development of a Clinical Decision Tool for Secondary or Tertiary Spine Care Specialists. PLoS ONE 9, 8 (August 2014), e104226. https://doi.org/10.1371/journal.pone.0104226
@@
- Venkatesh et al. (2016) Viswanath Venkatesh, James Thong, Hong Kong University of Science and Technology, Xin Xu, and The Hong Kong Polytechnic University. 2016. Unified Theory of Acceptance and Use of Technology: A Synthesis and the Road Ahead. Journal of the Association for Information Systems 17, 5 (May 2016), 328–376. https://doi.org/10.17705/1jais.00428
@@
Appendix A Sample Questions
A.1. The Socratic Question Taxonomy
Element of Thought | Sample Questions |
Purpose | What are we trying to accomplish? What is the purpose of …? |
A.2. The XAI Question Bank
Element | Sample Questions |
A.3. Our Taxonomy applied to Healthcare
ID | Element for Reflection | Sample Questions |
A.4. Our Taxonomy applied to Education
ID | Element for Reflection | Sample Questions |

### v3

@@ 5. Evaluation
Overall, both recognised the relevance of our taxonomy and the use of questions. They were confident that questions could help to re-evaluate initial assumptions of the physician, and counteract habitual priming, e.g., seeing similar yet different patient cases throughout the day. Importantly, as the surgeon noted, this priming also happens through the information presented by a DSS, which questions could help to mitigate. Moreover, they mentioned that the questions derived from the taxonomy could help to treat different patients more objectively by reducing discrepancies between cases, such as different patients describing their pain differently, i.e., pain catastrophising. As one remarked, given that many question types from the taxonomy are derived from data, it is difficult to ask about data points that are not contained in the dataset (Q3).111In this case, general questions like “Were all relevant information considered?” or “Did you consider factor x which is known to influence the outcome, but is not part of the data set” could help. To derive relevant questions, they suggested to talk to experts and ask for their reasoning process, as the usefulness of specific questions, as they noted, depend on the specific use case. Accordingly, they also pointed out that questions should help and not irritate.
5.1. Transferability of Our Taxonomy
To assess the applicability of our taxonomy in another domain, we discussed the finalised version with two postdoctoral researchers in the field of education and technology, and one doctoral researcher who is also a primary school teacher. These individual and informal conversations, each lasting about 30 minutes, were neither recorded nor transcribed. The aim of these conversations was to clarify whether the taxonomy is relevant for the educational domain, and to familiarise ourselves with some exemplary use cases and questions.
@@ 5.1. Transferability of Our Taxonomy
The three researchers shared two examples of so-called learning analytics dashboards (LAD), which educators can use to plan lessons, define learning objectives, track the learning progress of students, create exercises accordingly, and grade students.222https://snappet.org/ and https://www.gynzy.com/en. These LADs also recommend exercises and learning objectives based on student performance that teachers can use. In view of this, one researcher commented, “Teachers feel the need to be more critical, but do not know how” as they wonder whether they use the LAD correctly. Accordingly, the researchers noted that questions could contribute to the interpretation of the student’s progress shown in the LAD, and help consider factors that go beyond data points, as these systems “do not know everything”. All three researchers therefore confirmed the expected relevance and applicability of our taxonomy for stimulating teachers’ critical reflection in making decisions about students and their futures using recommendations of LADs.
@@
After the three feedback sessions, we created sample questions for the educational domain (Table D). We sent these to the three researchers and two of them added some questions and comments that helped us refine our derived questions. We then presented the sample questions, including our taxonomy, in a lab session attended by 12 people, from doctoral researchers to full professors. The research group, which explores the impacts of technologies on learning and teaching practices, recognised the usefulness of our taxonomy. As one participant commented, the sample questions we created are very close to the questions teachers already ask themselves. Nevertheless, the same researcher continued, our taxonomy could be informative to see which questions are covered and which are left out. In addition, another researcher noted that it is very promising that our taxonomy aligns with the questions that teachers ask in relation to algorithmic learning goal recommendations. Furthermore, one attendee noted that students doing the exercises in a LAD could also benefit from questions for critical reflection, such as considering the benefit of doing an exercise that might go against (short-term) preferences.
@@
In comparison to the medical sample questions (Table C), the wording of the educational sample questions differs. The structure and elements of critical reflection however remain the same for both. We therefore maintain that our taxonomy is transferable and can be applied to other decision-making domains without adaptation as it builds on general concepts of critical thinking.
@@ 5.2. Effectiveness of Questions
Another type of evaluation, beyond the scope of this paper, is to assess the effectiveness of the questions created based on the taxonomy in decision-making, and the extent to which they promote critical reflection and cognitive engagement of the decision-maker. Existing studies, primarily from the educational domain, point to the positive effects of Socratic questioning strategies on moral reasoning in nursing (Torabizadeh et al., 2018) and critical thinking skills during (healthcare) education of students (Ishtiaq et al., 2024; Jameel et al., 2025; Yang et al., 2005; Makhene, 2019; Ho et al., 2023). These studies demonstrate the validity and robustness of the Socratic questioning approach on which our proposed taxonomy is based, which makes the prospects for our taxonomy promising.
Nevertheless, assessing the effect of questions on decision-making remains an open challenge. Current evaluation methods for decision-making and overreliance on DSS typically only focus on the outcome of a decision, i.e., decision accuracy (Guo et al., 2024; Klingbeil et al., 2024). To measure reflection and cognitive engagement, however, it is important to consider the entire decision-making process, including the reasoning process, whether alternatives were considered, the decision-maker’s confidence in the final decision, and satisfaction with the outcome. Existing methods for measuring cognitive engagement are associated with considerable effort and technical barriers, such as electroencephalography (EEG) (Hassib et al., 2017; Berka et al., 2007), or the measurement of pupil dilation (Van Der Wel and Van Steenbergen, 2018). Consequently, an accessible evaluation framework would need to be developed in future work, e.g., a self-report scale that takes the entire decision-making process into account. For this, inspiration can be drawn from existing scales for factors associated with cognitive engagement, such as self-efficacy, motivation, and confidence (Lee et al., 2025; Hart et al., 2014; Greene, 2015), as well as from methods in the educational domain for assessing student’s cognitive engagement with technology (Vongkulluksn et al., 2022).
@@ 6. Discussion
We have presented a taxonomy of question types to support the systematic creation of questions that stimulate critical reflection in machine-assisted decision-making. As Figure 1 shows, the questions identified in the XAI question bank for human-centred XAI are very model- and data-centric. As a result, our taxonomy is more data-oriented. Nevertheless, our taxonomy can help to take into account human factors such as assumptions or cognitive biases of the decision-maker and stakeholder preferences as well as consequences of a decision to overcome the current algorithmic focus of XAI (section 2.2). In doing so, questions can stimulate discussion about the machine recommendation and its integration into the broader decision-making process.
@@ 6.1. Different Stakeholders
Our taxonomy is primarily targeted at decision-makers who operate DSS to make informed decisions. Accordingly, the aim of our taxonomy is not to formulate questions for the decision-maker to assess the performance of the DSS as this would unfairly shift the responsibility from developers and regulators to operators. The deployed models must be evaluated and tested beforehand, which requires a shift from individual human oversight to institutional oversight (Green, 2022). Considering this, other stakeholders could also benefit from critical reflection and thus our taxonomy. Developers constructing a model, for example, might ask themselves what assumptions they are making and whether the dataset used is suitable for solving a given problem. Auditors assessing compliance could benefit from our taxonomy when defining objectives for an audit of the model or dataset (Raji et al., 2020; Birhane et al., 2024). Moreover, the taxonomy could be used as a form of supporting AI literacy to help people remain more critical of machine-generated information, especially in view of misinformation (Zavolokina et al., 2024). Patients could be supported in identifying (treatment) preferences through reflection, thereby asking relevant questions to their healthcare providers (operating DSS), or being critical while interacting with a medical chatbot themselves. Students could be taught to question the relevance of retrieved information and could be helped to remain cognitively engaged in their technology-aided learning. Our proposed taxonomy could therefore benefit various stakeholders involved in the design, development, testing, and use of DSS or machine learning systems.
@@ 6.2. Creation of Questions
To facilitate the formulation of questions, we suggested that questions should address the levels of critical thinking according to Bloom’s taxonomy, i.e., evaluating, analysing, and creating. As such, it might be that questions cannot be answered directly and instead lead to further questions (Paul, 2019). To motivate further inquiry and engagement of the decision-maker, a brief explanation of why a question should be entertained could be given, as mentioned by the orthopaedic surgeon we spoke to.
It must be considered that questions can be leading and possibly steer the decision-maker towards an undesirable decision. Hence, we advise against the use of large language models, as the generated questions could be irrelevant and unreliable (Weidinger et al., 2022). Instead, researchers or designers ideally work together with experts to create useful questions. When creating the sample questions for the educational domain we realised that it is difficult to do so without having much knowledge about the respective use case, access to data, or the functionality of the DSS or learning analytics dashboard (LAD). Nevertheless, uncovering these blind spots might enable further inquiry into how the system works, and open up a dialogue between developers or service providers and domain experts or affected organisations, such as hospitals or schools.
Furthermore, certain questions might fall into multiple categories of our taxonomy, as different question types can overlap or build on other categories. The causal structure of a recommendation (Q4), for example, can entail questions about the data (Q1&Q3). We believe, however, that this is not a problem as it shows how different elements are connected to each other and it could rather help to formulate questions more precisely so that they can be delineated.
@@ 6.3. Presentation of Questions
To present the questions to decision-makers, one of the educational researchers we spoke to suggested that a poster of our taxonomy with sample questions could be displayed in the staff room to sensitise teachers to the different elements of the machine-assisted decision-making process. Alongside an analogue form and building on the above-mentioned proposals of DSS supporting deliberation and reflection (section 2.3), our taxonomy could inform the design of a technical implementation. As such, we introduced and envision a reflection machine (RM) that functions as an additional system or component of a DSS (Haselager et al., 2023; Cornelissen et al., 2022; Fischer, 2025). Based on input data, the recommendations, the functioning of the DSS derived from XAI, and the use of our taxonomy relevant questions can be created and presented to the decision-maker accordingly.
Supporting the decision-making process with questions might be an unfamiliar interaction. The design of the interface and the integration into existing workflows therefore influence the acceptance and perceived usefulness of a specific prototype (Lund, 2001), and thus the effectiveness of questions on cognitive engagement.
@@ 6.4. Cognitive Load
While future work needs to evaluate the effectiveness of questions, questions could interrupt the workflow and increase the cognitive load of the decision-maker. In the aforementioned study, in which explanations were presented at different times as a form of cognitive intervention, participants reported increased cognitive load, which is why they favoured these interventions less (Buçinca et al., 2021). Ultimately, however, the cognitive burden is unavoidable and inherent to critical reflection, as Dewey (1933, p.17) notes:
[…] yet thinking need not be reflective. For the person may not be sufficiently critical about the ideas that occur to him. He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled. One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching.
Nevertheless, the right balance between effort and effectiveness must be found. Questions must be well integrated into decision-making processes in order to enable the operator to constructively take the posed question into account. The competences, decision-making style, and need for cognition of the decision-maker play a role here.
@@ 7. Conclusion
Our taxonomy of 10 question types shifts the emphasis from machine recommendations and explanations to questions and human reflection. Careful selection and presentation of questions has the potential to increase cognitive engagement, mitigate overreliance on DSS and thus support the effective human oversight required by the European AI Act.
Appendix A The Socratic Question Taxonomy
| Element of Thought | Sample Questions |
| Purpose | What are we trying to accomplish? What is the purpose of …? |
Appendix B The XAI Question Bank
| Element | Sample Questions |
Appendix C Our Taxonomy Applied to Healthcare
| ID | Element for Reflection | Sample Questions |
Appendix D Our Taxonomy Applied to Education
| ID | Element for Reflection | Sample Questions |
@@ Acknowledgements.
We thank Sebastiaan van de Groes, at Radboud UMC, and Miranda van Hooff, at Sint Maartenskliniek and Radboud UMC, for their input on the construction of the taxonomy. Many thanks to Josh Ring and the Adaptive Learning Lab, in particular to Susan Verbunt-Janssen, Susanne de Mooij, and Zowi Vermeire for their valuable insights, and helping with the sample questions for the educational domain. Thanks to Cătălina Rățală and the Decision Science Group for their questions and feedback on a presentation of our paper. Comments from three ACM FAccT reviewers on a previous version, as well as comments from three AAAI/ACM AIES reviewers helped us to improve the paper. This research is funded by the Donders Centre for Cognition.References
@@
- Ely (2000) J. W Ely. 2000. A Taxonomy of Generic Clinical Questions: Classification Study. BMJ 321, 7258 (Aug. 2000), 429–432. https://doi.org/10.1136/bmj.321.7258.429
@@
- Fischer (2025) Simon Ws Fischer. 2025. A Reflection Machine to Support Critical Reflection During Decision-Making. In Companion Proceedings of the 30th International Conference on Intelligent User Interfaces. ACM, Cagliari Italy, 199–201. https://doi.org/10.1145/3708557.3716151
@@
- Ma et al. (2025) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma. 2025. Towards Human-AI Deliberation: Design and Evaluation of LLM-Empowered Deliberative AI for AI-Assisted Decision-Making. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–23. https://doi.org/10.1145/3706598.3713423
- Ma et al. (2024) Shuai Ma, Chenyi Zhang, Xinru Wang, Xiaojuan Ma, and Ming Yin. 2024. Beyond Recommender: An Exploratory Study of the Effects of Different AI Roles in AI-Assisted Decision Making. arXiv:2403.01791 [cs]
@@
- Miller (2023) Tim Miller. 2023. Explainable AI Is Dead, Long Live Explainable AI!: Hypothesis-driven Decision Support Using Evaluative AI. In 2023 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’23). ACM, Chicago IL USA, 333–342. https://doi.org/10.1145/3593013.3594001